what is the geometric mean of 8 and 32 ?
is it 16?

Answer by checkley77(12844)

Geometric mean of 8 and 32
L=a*r^(n-1)
L=32
a=8
n=3
32=8*r^(3-1)
32=8*r^2
r^2=32/8
r^2=4
r=sqrt4
r=+-2 ans.
Proof:
8,[16],32
8,[-16],32